OpenStudy (anonymous):

Could you explain to me the significance of vf'=[(m1-2m)/(m1+M2)]vi

OpenStudy (anonymous):

@arindameducationusc

OpenStudy (anonymous):

I am aware that this equations concerns with a collision between two objects.

OpenStudy (anonymous):

Whether this equation concerns with linear collision or two dimensional collision I don't know.

OpenStudy (arindameducationusc):

yes?

OpenStudy (arindameducationusc):

can you please explain the terms? what are they?

OpenStudy (arindameducationusc):

ok ok wait....

OpenStudy (anonymous):

Sorry I was in the washroom

OpenStudy (anonymous):

Alright so this concerns with conservation of momentum.

OpenStudy (anonymous):

I am unsure as to what v' flick means

OpenStudy (arindameducationusc):

hey,,,, its one dimention..... As component is not there... I haven't revised that, but just went through my notes roughly

OpenStudy (arindameducationusc):

Its a good thing you reminded me, Once I revise NLM and Work, Energy, Power, I will go through Collisions pretty thoroughly. but ya, i am pretty sure, this is one dimention

OpenStudy (arindameducationusc):

For 2D The equation is (don't ask me how as I haven't revised yet) is=> \[m _{1}u _{1}\sin \alpha+m _{2}u _{2}\sin \beta=m _{1}v _{1}\sin \theta _{1}+m _{2}v _{2}\sin \theta _{2}\]

OpenStudy (arindameducationusc):